# Advances in Nonmonotone Proximal Gradient Methods merely with Local Lipschitz Assumptions in the Presense of Kurdyka–Łojasiewicz Property: A Study of Average and Max Line Search

### v1

@@
Abstract. The proximal gradient method is a standard approach to solve the composite minimization problems where the objective function is the sum of a continuously differentiable function and a lower semicontinuous, extended-valued function. For both monotone and nonmonotone proximal gradient methods, the convergence theory has traditionally replied heavily on the assumption of global Lipschitz continuity. Recent works have shown that the monotone proximal gradient method, even when the local Lipschitz continuity (rather than global) is assumed, converges to the stationarity globally in the presence of Kurdyka–Łojasiewicz Property. However, how to extend these results from monotone proximal gradient method to nonmonotone proximal gradient method (NPG) remains an open question. In this manuscript, we consider two types of NPG: those combined with average line search and max line search, respectively. By partitioning of indices into two subsets, one of them aims to achieve a decrease in the functional sequence, we establish the global convergence and rate-of-convergence (same as the monotone version) results under the KL property, merely requiring the local Lipschitz assumption, and without an a priori knowledge of the iterative sequence being bounded. When our work is almost done, we noticed that [17] presented the analogous results for the NPG with average line search, whose partitioning of index set is totally different with ours. Drawing upon the findings in this manuscript and [17], we confidently conclude that the convergence theory of NPG is independent on the specific partitioning of the index set.
Keywords. Composite problems $\cdot$ Nonmonotone proximal gradient methods $\cdot$ Kurdyka-Łojasiewicz property $\cdot$ Local Lipschitz continuity
@@ 1 Introduction
where $\mathbb{X}$ is an Euclidean space, i.e., real and finite-dimensional Hilbert space, $f:\mathbb{X}\to\mathbb{R}$ is continuously differentiable, $g:\mathbb{X}\to\overline{\mathbb{R}}$ is assumed to be merely lower semicontinuous. We claim that the unconstrained composite optimization problem (Q) is in a totally nonconvex setting. This type of optimization problems have a wide range of applications in practice, like machine learning, image processing, and data science [4, 11, 8].
@@
In this manuscript, we are interested in two classes of nonmonotone proximal gradient method to solve (Q), where the nonmonotone line search technique based on monitoring the weighted average of the objective value at all iterates by Zhang and Hager [28], and the maximum objective value attained from the latest iterates by Grippo et al. [13], respectively. More precisely, some iterate $x^{k+1}\in\operatorname{dom}q$ is acceptable provided that the corresponding objective $q(x^{k+1})$ is less than the merit
$\Phi_{k}:=(1-p)\Phi_{k-1}+pq(x^{k})\qquad\text{and}\qquad q(x^{l(k)}):=\max_{j=0,\ldots,\min\{m,k\}}q(x^{k-j})$
@@
for the average and max taste, respectively. Note that $m\in\mathbb{N}$ and $p\in(0,1]$ (usually) are constants in order to control the level of nonmonotonicity, i.e., for the smaller $p$, or the larger $m$, the corresponding metric function can attain larger values, hence imposing a weaker condition for acceptance of $x^{k+1}$ compared to the monotone proximal gradient methods [12]. The values of $p=1$ and $m=0$ lead to the monotone behavior, i.e., the nonmotone proximal gradient method degenerates into the monotone one, please see [18, Algorithm 3.1] and [15, Algorithm 3.1].
@@
The convergence theory associated with (monotone and nonmonotone) proximal gradient methods mostly require the derivative of the smooth part of the objective function to be global Lipschitz continuous, which, however, is quite restrictive assumption in practical scenarios. [15, Examples 3.6 and 3.7], as two classical examples, revealed that the global Lipschitz assumption on the gradient of $f$ is typically violated, whereas the local Lipschitz condition is often satisfied. [18] readdressed the classical monotone proximal gradient method [18, Algorithm 3.1] and NPG with max line search [18, Algorithm 4.1], and provided (subsequential) convergence results only with local Lipschitz assumption (not global any more). As for the convergence of the whole sequence generated by the proximal gradient method with local Lipschitz assumption, [15] is the first one to reveal this mystery for monotone proximal gradient in the presence of the Kurdyka-Łojasiewicz property of the objective function. [12] considered an adaptive nonmonotone proximal gradient scheme based on an averaged metric function and established (subsequential) convergence results with local Lipschitz assumption of the gradient of $f$, where the global worst-case rates for the iterates and a (subsequential) stationarity measure were also derived. Note that the numerical behaviours and competitiveness has already been investigated in many literature before, for example see [26, 12, 25, 20, 21, 27], this manuscript focuses more on theoretical aspects of the method (without numerical experiments), showing qualitative and even quantitative properties of the NPG in a more general framework which allows for applications e.g. in the context of multiplier-penalty methods [16]. More precisely, we, in this manuscript, under the same assumptions, consider the convergence theory of the entire sequence generated by NPG with both average and max backtracking line search when KL property being employed.
@@
Note that the corresponding convergence results about NPG with max line search has been analysed in [25], where an auxiliary sequence is necessarily prerequisite to be convergent, which, however, is not needed at all in our technique, moreover, where the boundedness of subdifferential of the objective function is needed for all iterates, which, to the best of our knowledge, fails to be satisfied under the local Lipschitz continuous assumption. However, in this manuscript, no any prerequisite is assumed and the corresponding convergence and rate-of-convergence results obtained in [25] can be also achieved merely by requiring the local Lipschitz continuity of $\nabla f$.
@@
The primary challenge posed by NPG, in fact, is that the generated functional sequence may not be decreasing, to date, whereas the (sufficient) descent property seems to be necessary for the technical proof when using KL property [2, 3, 7, 8]. To the best of our knowledge, nearby all existing studies addressing this issue reply on the partition of the index sets [25, 17, 26]. Specifically, these works divided the indices into two subsets: one subset ensures a decrease in the functional sequence, while the complement accounts for the remaining cases. The recent paper [17] considered the convergence results of the NPG with average line search under the KL property, merely requiring the local Lipschitz continuity of $\nabla f$, where the index sets are defined as
$K_{1}:=\{k\in\mathbb{N}\,|\,q(x^{k})\leq\Phi_{k+m}\},\quad K_{2}:=\{k\in\mathbb{N}\,|\,q(x^{k})>\Phi_{k+m}\},$
@@
where $m$ is some predetermined integer relying on $p_{\min}$ to facilitate the convergence analysis. This partitioning originates from [26] and the subsequent convergence analysis heavily depend on the techniques therein, where the key component is exploiting the quantitative relationship between $\sum_{i=k}^{k+m-1}\sqrt{\Phi_{i-1}-\Phi_{i}}$ and $\sqrt{\Phi_{i-1}-\Phi_{i}}$, see [17, Lemma 4.4] and [26, Theorem 2.1]. To be honest, we and [17] happened to consider the same topic almost simultaneously. In this manuscript, for the NPG with average line search, to the best of our knowledge, we are the first one to propose a different partitioning of the index sets from [26]:
@@
where $\mu\in(0,\frac{1}{2}\delta{p_{\min}\gamma_{\min}}]$ is a constant, and $p_{\min}$ and $\gamma_{\min}$ are parameters derived from the algorithm. Actually, different partitioning of the index sets necessitates the different techniques for proof. However, our approach yields more natural and concise proofs compared to [17], as it eliminates the need for the mentioned-above quantitative relationship ([17, Lemma 4.4] and [26, Theorem 2.1]). Our analysis reproduces the corresponding convergence and rate-of-convergence presented in [17]. Furthermore, [17] leaves the theoretical analysis of NPG methods with max line search for future work, we provide a comprehensive analysis in this manuscript. The decision to present this part of work (NPG with average line search), despite its similarities to [17], is also justified by a noteworthy finding: both works demonstrate that the convergence results of nonmonotone proximal gradient methods are independent on the specific partitioning of the index sets. This insight contributes a significant advancement to the understanding of this method.
This manuscript is organized as follows: We first recall some background knowledge in Section 2. We then present the basic properties and convergence results of the NPG with average line search and with max line search in Section 3 and Section 4, respectively. Section 5 concludes this manuscript.
@@ 2 Preliminaries
denotes the distance of the point $x$ to the set $A$ with $\operatorname{dist}(x,\emptyset):=\infty$. For given $\varepsilon>0$, $B_{\varepsilon}(x):=\{y\in\mathbb{X}\,|\,\left\|y-x\right\|\leq\varepsilon\}$ denotes the closed $\varepsilon$-ball around $x$.
The continuous linear operator $f^{\prime}(x)\colon\mathbb{X}\to\mathbb{R}$ denotes the derivative of the continuously differentiable function $f\colon\mathbb{X}\to\mathbb{R}$ at $x\in\mathbb{X}$, and we will make use of $\nabla f(x):=f^{\prime}(x)^{*}1$ where $f^{\prime}(x)^{*}\colon\mathbb{R}\to\mathbb{X}$ is the adjoint of $f^{\prime}(x)$. This way, $\nabla f$ is a mapping from $\mathbb{X}$ to $\mathbb{X}$.
We further say that a sequence $\{x^{k}\}\subset\mathbb{X}$ converges Q-linearly to $x^{*}\in\mathbb{X}$ if there is a constant $c\in(0,1)$ such that the inequality
@@
holds for all sufficiently large $k\in\mathbb{N}$. Furthermore, $\{x^{k}\}$ is said to converge R-linearly to $x^{*}$ if we have
@@
Note that this R-linear convergence holds if there exist constants $\omega>0$ and $\mu\in(0,1)$ such that $\|x^{k}-x^{*}\|\leq\omega\mu^{k}$ holds for all sufficiently large $k\in\mathbb{N}$, i.e., if the expression $\|x^{k}-x^{*}\|$ is dominated by a Q-linearly convergent null sequence.
Let us fix a merely lower semicontinuous function $q\colon\mathbb{X}\to\overline{\mathbb{R}}$ and pick $x\in\operatorname{dom}q$ where $\operatorname{dom}q:=\{x\in\mathbb{X}\,|\,q(x)<\infty\}$ denotes the domain of $q$. Then the set
@@
is called the regular (or Fréchet) subdifferential of $q$ at $x$. Furthermore, the set
@@
is well known as the limiting (or Mordukhovich) subdifferential of $q$ at $x$. Clearly, we always have ^$\widehat{\partial}q(x)\subset\partial q(x)$ by construction of these sets. Whenever $q$ is convex, both subdifferentials coincide, i.e.,
^$\widehat{\partial}q(x)=\partial q(x)=\{\eta\in\mathbb{X}\,|\,\forall y\in\operatorname{dom}q\colon\,q(y)\geq q(x)+\langle\eta,y-x\rangle\}$
is valid in this situation. By definition of the regular subdifferential, it is clear that whenever $x^{*}\in\operatorname{dom}q$ is a local minimizer of $q$, then $0\in\widehat{\partial}q(x^{*})$ hold by Fermat’s rule [22, Proposition 1.30(i)]. Thus, the inclusion $0\in\partial q(x^{*})$ is a necessary optimality condition for $x^{*}$ being a local minimizer of $q$ as well.
@@
holds, see [22, Proposition 1.30(ii)]. Fermat’s rule shows that the optimality condition
@@
holds at any local minimizer $x^{*}\in\operatorname{dom}\phi$ of the composite optimization problem (Q). Any point $x^{*}\in\operatorname{dom}\phi$ satisfying this necessary optimality condition will be called an M-stationary point of (Q).
We next introduce the famous Kurdyka–Łojasiewicz property which plays a central role in this manuscript. The version of this property stated below is a generalization of the classical Kurdyka–Łojasiewicz inequality for nonsmooth functions as introduced in [2, 5, 6] and afterwards used in the local convergence analysis of several nonsmooth optimization methods, see [1, 3, 7, 9, 10, 23, 24] for a couple of examples.
@@
Let $g\colon\mathbb{X}\to\overline{\mathbb{R}}$ be lower semicontinuous. We say that $g$ has the KL property, where KL abbreviates Kurdyka–Łojasiewicz, at $x^{*}\in\{x\in\mathbb{X}\,|\,\partial g(x)\neq\emptyset\}$ if there exist a constant $\eta>0$, a neighborhood $U\subset\mathbb{X}$ of $x^{*}$, and a continuous concave function $\chi\colon[0,\eta]\to[0,\infty)$ which is continuously differentiable on $(0,\eta)$ and satisfies $\chi(0)=0$ as well as $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$ such that the so-called KL inequality
@@
holds for all $x\in U\cap\big{\{}x\in\mathbb{X}\,|\,g(x^{*})<g(x)<g(x^{*})+\eta\big{\}}$. The function $\chi$ from above is referred to as the desingularization function.
@@
3 NPG with Average Line Search and Convergence Results
We first propose the nonmonotone proximal gradient method with average line search in Algorithm 3.1.
Algorithm 3.1 Nonmonotone Proximal Gradient Method with Average Line Search

### v3

# Convergence analysis of nonmonotone proximal gradient methods under local Lipschitz continuity and Kurdyka–Łojasiewicz property

@@
Abstract. The proximal gradient method is a standard approach for solving composite minimization problems in which the objective function is the sum of a continuously differentiable function and a lower semicontinuous, extended-valued function. The traditional convergence theory for both monotone and nonmonotone variants replies heavily on the assumption of global Lipschitz continuity of the gradient of the smooth part of the objective function. Recent work has shown that monotone proximal gradient methods converge globally only when the local (rather than global) Lipschitz continuity is assumed, provided that the Kurdyka–Łojasiewicz (KL) property holds. However, these results have not been extended to nonmonotone proximal gradient (NPG) methods. In this manuscript, we consider two types of NPG methods: those combined with the average line search and the max line search, respectively. By partitioning indices into two subsets, one of which aims to achieve a sufficient decrease in the functional sequence, we establish global convergence and rate-of-convergence results using the local Lipschitz continuity and the KL property, without requiring boundedness of the iterates. While finalizing this work, we noticed that [18] presented analogous results for the NPG method with average line search, but with a different partitioning strategy. Together, we confidently conclude that the convergence theory of the NPG method is independent on index partitioning choices.
Keywords. Composite problems $\cdot$ Nonmonotone proximal gradient methods $\cdot$ Average and max line searches $\cdot$ Kurdyka-Łojasiewicz property $\cdot$ Local Lipschitz continuity
@@ 1 Introduction
where $\mathbb{X}$ is an Euclidean space, i.e., real and finite-dimensional Hilbert space, $f:\mathbb{X}\to\mathbb{R}$ is continuously differentiable, $g:\mathbb{X}\to\overline{\mathbb{R}}$ is assumed to be merely lower semicontinuous. Note that (Q) is totally nonconvex, which has a wide range of applications in practice, like machine learning, image processing, and data science [4, 12, 8].
@@
In this manuscript, we are interested in two types of nonmonotone proximal gradient methods for solving (Q), where the nonmonotone line search techniques are based on different nonmonotone criteria. Specifically, we consider the approach of Zhang and Hager [29], which monitors the weighted average of the objective values over the past iterates, and the approach of Grippo et al. [14], which uses the maximum objective value attained among the most recent iterates. In both cases, a candidate iterate $x^{k+1}\in\operatorname{dom}q$ is acceptable if the corresponding objective $q(x^{k+1})$ is less than a designated merit function, defined as
[…]
@@
where $m\in\mathbb{N}$ and $p\in(0,1]$ are constants, $l(k)\in\{k-m,\ldots,k\}$. Note that $p$ and $m$ aim to control the level of nonmonotonicity. For the smaller $p$, or the larger $m$, the corresponding metric function attains larger values, hence imposing a weaker condition for the acceptance of $x^{k+1}$ compared to the monotone proximal gradient methods [13]. The values of $p=1$ and $m=0$ lead to the monotone behavior, i.e., the nonmotone proximal gradient method degenerates into the monotone one, please see [19, Algorithm 3.1] and [16, Algorithm 3.1].
@@
The convergence theory associated with (monotone and nonmonotone) proximal gradient methods primarily requires the derivative of the smooth part of the objective function to be globally Lipschitz continuous. However, the requirement is quite restrictive in practical scenarios. [16, Examples 3.6 and 3.7], as two classical examples, revealed that the global Lipschitz assumption on $\nabla f$ is typically violated, whereas the local Lipschitz condition is often satisfied. [19] readdressed the classical monotone proximal gradient method [19, Algorithm 3.1] and the NPG method with a max line search [19, Algorithm 4.1], and provided subsequential convergence results based solely on the local (not global any more) Lipschitz assumption. Regarding the convergence of the entire sequence generated by proximal gradient methods using the local Lipschitz assumption, [16, 9] revealed this mystery for monotone proximal gradient methods in the presence of the Kurdyka-Łojasiewicz (KL) property of the objective function. [13] considered an adaptive nonmonotone proximal gradient scheme based on an averaged metric function and established the subsequential convergence results with the local Lipschitz assumption of $\nabla f$, where the global worst-case rates for the iterates and the subsequential stationarity were also derived. Note that numerical behaviours and competitiveness have already been investigated in many literature, for example see [27, 13, 26, 21, 22, 28]. In this manuscript, we emphasize theoretical aspects of the method (without numerical experiments), showing qualitative and even quantitative properties of NPG methods with a more general framework, facilitating various applications, such as in the context of multiplier-penalty methods [17]. More precisely, we, in this manuscript, consider the convergence theory of the entire sequence generated by the NPG methods with both average and max backtracking line search techniques under the same assumptions when the KL property is employed.
@@
The corresponding convergence results for the NPG method with max line search have been analyzed in [26], in which the presence of an auxiliary sequence is a necessary prerequisite for convergence. However, such requirement is not required in our technique. One the other hand, the boundedness of subdifferential of the objective function is needed in [26] for all iterates, which, to the best of our knowledge, does not hold under the local Lipschitz continuity. In this manuscript, note that no any prerequisite is assumed and the corresponding convergence and rate-of-convergence results obtained in [26] can be also achieved by merely requiring the local Lipschitz continuity of $\nabla f$.
@@
The primary challenge posed by the NPG method, in fact, is that the generated functional sequence may not exhibit a sufficient decrease. To date, the sufficient descent property appears to be necessary for the technical proof when the KL property is employed [2, 3, 7, 8]. To the best of our knowledge, nearby all existing studies addressing this issue reply on partitioning the index sets [26, 18, 27]. Specifically, these works divide the indices into two subsets: one subset ensures a sufficient decrease in the functional sequence, while the complement accounts for the remaining case. A recent paper [18] considered the convergence results of the NPG method with average line search under the KL property, merely requiring the local Lipschitz continuity of $\nabla f$, where the index sets are defined as follows:
$K_{1}:=\{k\in\mathbb{N}\,|\,q(x^{k})\leq\Phi_{k+\bar{m}}\},\quad K_{2}:=\{k\in\mathbb{N}\,|\,q(x^{k})>\Phi_{k+\bar{m}}\},$
@@
where $\bar{m}$ is a predetermined integer that relies on $p_{\min}$ to facilitate the convergence analysis. This partitioning originates from [27] and the subsequent convergence analysis heavily depends on the techniques presented therein, where the key component is exploiting the quantitative relationship between $\sum_{i=k}^{k+\bar{m}-1}\sqrt{\Phi_{i-1}-\Phi_{i}}$ and $\sqrt{\Phi_{i-1}-\Phi_{i}}$, see [18, Lemma 4.4] and [27, Theorem 2.1]. To be honest, we and [18] happened to consider the same topic almost simultaneously. In this manuscript, for the NPG method with average line search, to the best of our knowledge, we are the first to propose a different partitioning of the index sets from [27, 18], as follows:
@@
where $\mu\in(0,\frac{1}{2}\delta{p_{\min}\gamma_{\min}}]$ is a constant, and $\delta$, $p_{\min}$, and $\gamma_{\min}$ are parameters derived from the algorithm. Actually, different partitionings of the index sets necessitate different techniques for proofs. However, our approach yields more natural and concise proofs than those in [18], because it eliminates the need for the mentioned-above quantitative relationship ([18, Lemma 4.4] and [27, Theorem 2.1]). Furthermore, [18] presents the theoretical analysis of the NPG method with a max line search for future work, we conduct the comprehensive analysis in this manuscript. The decision to present this part of work (the NPG method with average line search), despite its similarities to [18], is also justified by a noteworthy finding: both works demonstrate that the convergence results of NPG methods are independent on the specific partitioning of the index sets. This insight significantly advances the understanding of the methods.
This manuscript is organized as follows: We first recall some background knowledge in Section 2. We then present the basic properties and convergence results of NPG methods with average line search and max line search in Section 3 and Section 4, respectively. Section 5 concludes this manuscript.
@@ 2 Preliminaries
as the distance of the point $x$ to the set $A$ with $\operatorname{dist}(x,\emptyset):=\infty$. For any given $\varepsilon>0$, denote $B_{\varepsilon}(x):=\{y\in\mathbb{X}\,|\,\left\|y-x\right\|\leq\varepsilon\}$ as the closed $\varepsilon$-ball centered at $x$.
Let the function $f\colon\mathbb{X}\to\mathbb{R}$ be continuously differentiable, the continuous linear operator $f^{\prime}(x)\colon\mathbb{X}\to\mathbb{R}$ denotes its derivative at $x\in\mathbb{X}$, and we use $\nabla f(x):=f^{\prime}(x)^{*}1$ where $f^{\prime}(x)^{*}\colon\mathbb{R}\to\mathbb{X}$ is the adjoint of $f^{\prime}(x)$. So, $\nabla f$ is a mapping from $\mathbb{X}$ to $\mathbb{X}$.
A sequence $\{x^{k}\}\subset\mathbb{X}$ is said to converge Q-linearly to $x^{*}\in\mathbb{X}$ if there is a constant $c\in(0,1)$ such that the inequality
@@
holds for all sufficiently large $k\in\mathbb{N}$. Furthermore, $\{x^{k}\}$ converges R-linearly to $x^{*}$ if and only if we have
@@
Note that the R-linear convergence holds if there exist constants $a>0$ and $b\in(0,1)$ such that $\|x^{k}-x^{*}\|\leq ab^{k}$ holds for all sufficiently large $k\in\mathbb{N}$, in other words, $\|x^{k}-x^{*}\|$ is dominated by a Q-linearly convergent null sequence.
Let a function $q\colon\mathbb{X}\to\overline{\mathbb{R}}$ be merely lower semicontinuous and let $x\in\operatorname{dom}q$ where $\operatorname{dom}q:=\{x\in\mathbb{X}\,|\,q(x)<\infty\}$ denotes the domain of $q$. Then
@@
is called the regular (or Fréchet) subdifferential of $q$ at $x$. Furthermore,
@@
is denoted as the limiting (or Mordukhovich) subdifferential of $q$ at $x$. Naturally, ^$\widehat{\partial}q(x)\subset\partial q(x)$ always holds. Whenever $q$ is convex, both subdifferentials coincide, i.e.,
^$\widehat{\partial}q(x)=\partial q(x)=\{\eta\in\mathbb{X}\,|\,\forall y\in\operatorname{dom}q\colon\,q(y)\geq q(x)+\langle\eta,y-x\rangle\}.$
By Fermat’s rule [23, Proposition 1.30(i)], whenever $x^{*}\in\operatorname{dom}q$ is a local minimizer of $q$, then $0\in\widehat{\partial}q(x^{*})$ hold, consequently $0\in\partial q(x^{*})$ holds as well.
@@
holds, see [23, Proposition 1.30(ii)]. Fermat’s rule shows that the optimality condition
@@
holds at any local minimizer $x^{*}\in\operatorname{dom}\phi$ of (Q). Any point $x^{*}\in\operatorname{dom}\phi$ satisfying this necessary optimality condition is called an M-stationary point of (Q).
We next introduce the Kurdyka–Łojasiewicz (KL) property which plays a vital role in this manuscript. The definition stated below is a generalization of the classical KL inequality for nonsmooth functions as introduced in [2, 5, 6] and afterwards used in the convergence analysis of many nonsmooth optimization algorithms, see [1, 3, 7, 10, 11, 24, 25] for a couple of examples.
@@
Let $g\colon\mathbb{X}\to\overline{\mathbb{R}}$ be lower semicontinuous. We say that $g$ has the KL property at $x^{*}\in\{x\in\mathbb{X}\,|\,\partial g(x)\neq\emptyset\}$ if there exist a constant $\eta>0$, a neighborhood $U\subset\mathbb{X}$ of $x^{*}$, and a continuous concave function $\chi\colon[0,\eta]\to[0,\infty)$ which is continuously differentiable on $(0,\eta)$ and satisfies $\chi(0)=0$ as well as $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$ such that the so-called KL inequality
@@
holds for all $x\in U\cap\big{\{}x\in\mathbb{X}\,|\,g(x^{*})<g(x)<g(x^{*})+\eta\big{\}}$. The function $\chi$ is denoted as the desingularization function.
@@
3 The NPG Method with Average Line Search and Convergence Results
We first recall the nonmonotone proximal gradient method with average line search in Algorithm 3.1.
Algorithm 3.1 A Nonmonotone Proximal Gradient Method with Average Line Search